Run 11 timeline

From RHIC Spin Group
Jump to navigationJump to search

Last modified by Dmitri Smirnov on 21-12-2011


Run 9 timeline Run 11 timeline Run 12 timeline


Feb 11, 2011 @ ~2200
Physics run declared. Fill 15154
Feb 16, 2011
The downstream polarimeters moved from the tunnel to the counting room
Feb 17, 2011
the jump quads was in use(Polarization profile should improve). Fill 15190. The same time, PHENIX rotator was on.
Feb 19, 2011
Downstream polarimeter Y1D start reading out in the counting room after timing adjustments
Feb 22, 2011
Downstream polarimeter B2D start reading out in the counting room after timing adjustments
Feb 23, 2011 @ 1800
Special APEX session. Fill 15221
Feb 24 2011
9MHz cavity in operation, Fill 15232 (short one), 15233.
Feb 25, 2011 @ ~1630
The analyzing power used in online analysis was scaled. Online polarization values scaled down by ~18%. Fill 15239.
Mar 6 - 18 2011
RHIC cryo problems, no beam
Mar 7, 2011
Y1D timing params adjusted to agree with Y2U, ~7 nS t shift overall
Mar 17, 2011
Y1D pulser timing adjusted to move away from banana, starting fill 15300. First fill with Y1D data is 15342
Mar 23, 2011
Lower vertical tune between 100-250GeV to 0.673(was 0.675), starting fill 15325.
Apr 7, 2011 @ ~1900 - Apr 8, 2011 @ ~0030
Ramp up/down study. Fills 15415, 15416, 15417, 15418
Apr 8, 2011 @ ~1800
No vertical targets left in Y1D (?)
Apr 8, 2011 @ 1900
STAR rotators on. Fill 15424.
Apr 11, 2011 @ 2036
Keith tuned the linac phase, smaller emittances afterwards. Fill 15438.
Apr 13, 2011 @ 0230
last store before 9MHz failure. Fill 15444.
Apr 13, 2011 @ 2336
first store after switching to 28MHz cavities, low intensity. Raised Qy between 100-250GeV to 0.675. Reduced chrom. in this region, too. Fill 15452.
Apr 15, 2011
Ramps with different Qy between 100-250GeV. Fills 15460 (0.72), 15461 (0.71), 15462 (.671;failed).
Apr 16, 2011 @ ~0300
The Hamamatsu detectors (90-degree) in B2D failed and were excluded from readout. The failure signature is like the famous problem we experienced in the tunnel.
Apr 18, 2011
Detector 4 in B2D became noisy?
Apr 18, 2011 @ ~1230 — Apr 19, 2011 @ ~1000
Special injection fills at 24 GeV beam energy for H-jet calibration. Fills 15473, 15474, and 15475. The last fill had only blue beam


Last modified by Dmitri Smirnov on 21-12-2011